Write the linear equation in slope-intercept form;: -3x-4y-19=0
Answer by tutorcecilia(2144)

y=mx+b
-3x-4y-19=0 [isolate the y-term]
-3x+3x-4y-19=0+3x
-4y-19=3x
-4y-19+19=3x+19
-4y=3x+19
-4y/-4=3x/-4+19/-4
y=-3/4x-19/4